Calculate the nominal annual cost of trade credit under each of the following terms. Assume a 365-day year. Do not round intermediate calculations. Round your answers to two decimal places.

  1. 1/15, net 30.

      %

  2. 2/10, net 55.

      %

  3. 3/10, net 50.

      %

  4. 2/10, net 50.

      %

  5. 2/15, net 40.

      %
Answer #1

Cost of trade credit: = (1+D=(1-D)(365:n)-1 In is Number of days after discount period a Nominal annual cost of trade credit

*Please rate thumbs up